# Full-waveform earthquake source inversion using simulation-based inference

### v1

@@
Earthquake source observations – Waveform inversion – Bayesian inference – Inverse theory – Machine learningThis paper presents a novel framework for full-waveform seismic source inversion using simulation-based inference (SBI). Traditional probabilistic approaches often rely on simplifying assumptions about data errors, which we show can lead to inaccurate uncertainty quantification. SBI addresses this limitation by building an empirical probabilistic model of the data errors using machine learning models, known as neural density estimators, which can then be integrated into the Bayesian inference framework. We apply the SBI framework to point-source moment tensor inversions as well as joint moment tensor and time-location inversions. We construct a range of synthetic examples to explore the quality of the SBI solutions, as well as to compare the SBI results with standard Gaussian likelihood-based Bayesian inversions. We then demonstrate that under real seismic noise, common Gaussian likelihood assumptions for treating full-waveform data yield overconfident posterior distributions that underestimate the moment tensor component uncertainties by up to a factor of 3. We contrast this with SBI, which produces well-calibrated posteriors that generally agree with the true seismic source parameters, and offers an order-of-magnitude reduction in the number of simulations required to perform inference compared to standard Monte Carlo techniques. Finally, we apply our methodology to a pair of moderate magnitude earthquakes in the North Atlantic. We utilise seismic waveforms recorded by the recent UPFLOW ocean bottom seismometer array as well as by regional land stations in the Azores, comparing full moment tensor and source-time location posteriors between SBI and a Gaussian likelihood approach. We find that our adaptation of SBI can be directly applied to real earthquake sources to efficiently produce high quality posterior distributions that significantly improve upon Gaussian likelihood approaches.
@@ 1 Introduction
However, there are still some key remaining challenges in seismic source inversions. For example, the study of earthquakes in mid-ocean settings is typically limited by large distances to seismic stations, which provide useful but restricted information about the source process (López-Comino et al., 2015). Local/regional data from marine seismology experiments can help address this issue, but such expensive experiments are usually limited to short time periods and the data collected are often noisy (e.g., Stähler et al., 2016; Tsekhmistrenko et al., 2024). Crucially, robust uncertainty quantification is still lacking in many earthquake source studies, which limits meaningful interpretations (Wéber, 2006). Over the past twenty years, the Bayesian treatment of seismic source inversion has received increasing interest (Lomax et al., 2000; Tarantola, 2005; Vackář et al., 2017; Vasyura-Bathke et al., 2020). Seismic source inversion can be complicated by inter-parameter tradeoffs and strong non-uniqueness (Kawakatsu, 1996; Dufumier & Rivera, 1997; Ford et al., 2010), and previous work has demonstrated that classical statistical analysis of standard errors can significantly underestimate uncertainty in source parameters (Duputel et al., 2012; Valentine & Trampert, 2012). A number of studies have demonstrated that thorough Bayesian treatment of seismic events uncovers uncertainties and parameter tradeoffs, often revealing greater complexity than simple maximum likelihood solutions (Duputel et al., 2012; Pugh et al., 2016; Hu et al., 2023). Simutė et al. (2023), for example, demonstrated how inferred focal mechanisms and their distributions are sensitive to the seismic frequencies considered in the inversion, and that a Bayesian analysis of seismic source inversion can produce slightly multimodal posteriors. In addition, recent work has demonstrated that full-waveform probabilistic inversions can be highly sensitive to the noise model parametrisation (Duputel et al., 2012; Mustać & Tkalčić, 2016; Mustać et al., 2018).
@@
To the best of our knowledge this study is the first application of SBI in seismology. We therefore begin this paper by presenting the motivation and prior work in SBI, introducing modern ML techniques for probability distribution modelling, and presenting a data compression technique to treat the full-waveform source inversion problem with SBI. We then present our methodology and apply it to a range of synthetic waveform inversions. Subsequently we use an ocean bottom seismology dataset recently collected in the Atlantic region (the UPFLOW dataset; Tsekhmistrenko et al. (2024)) to study two oceanic earthquakes. This allows us to probe the advantages and pitfalls of using SBI in earthquake source inversions using noisy seismic data.
@@ 3.3 Neural density estimators (NDEs)
The task of inference then relies on finding and training a model capable of modelling complex probability distributions. Much of the success of SBI hinges on the advancements made in the flexibility and efficiency of NDEs over the last decade. Neural density estimation is a conceptually simple task at its core; given a dataset of observations of , a NDE must be trained to estimate the density across . Conditional probabilities can be similarly treated, with a dataset of allowing the empirical modelling of e.g. density . In practice, the challenge in performing this modelling is to design a flexible density estimator with a tractable (ideally easy to compute) density function.
@@
where is the Jacobian of T, i.e., the matrix of partial derivatives with respect to components of . One type of normalising flow are autoregressive models, which proceed by learning to model a succession of conditional distributions, thereby constructing the full probability distribution by applying the chain rule:
@@
These flows also have the useful property that they can be composed together, in effect learning to parametrise successive transformations, allowing for significant improvements in the flexibility of the model. MAFs, which we utilise in this work, are constructed by stacking together several MADE networks and varying the autoregression order from Eq. 5 between each MADE. In this work, MAFs are trained to model the posterior , so each transformation parametrised by a MADE is also conditioned on the observation . This architecture is depicted in Fig. 2.
Once the network architecture has been selected, the NDE can be trained to maximise the posterior density over a set of observations , by minimising the objective function . This straightforward formulation can be combined with standard stochastic gradient descent techniques common to most ML applications to train the NDE .
@@ 3.5 Optimal Score Compression
where the Fisher information matrix is defined as for log-likelihood , and . Here, represents an expectation over the model parameters. Eq. 7 states that the variance of the statistic over the model parameters can be bounded from below (where equality holds for an “optimal” statistic). Stated informally, the aim is then to minimise the variance over for a given set of model parameters , hence maximally constraining when inferring it from an observation .
@@
where is the synthetic data at the expansion point , and is the gradient of this data vector with respect to the model parameters. Eq. 10 thus defines the compression operator for the optimal score compression scheme. This form of has a familiar interpretation in the context of seismic inverse theory; it is the vector of sensitivity kernel coefficients for each of the model parameters of interest (i.e., coefficients of the first partial derivatives of the data with respect to the model parameters), computed with reference to a fixed expansion point .
@@
where the Fisher information for the Gaussian likelihood in Eq. 8. This form of the summary statistic corresponds to a local maximum likelihood estimate (MLE) of the model parameters, , given the observation ; equivalently, it is an estimate of the model parameters after a single step of least-squares from expansion point . Alsing & Wandelt (2018) also showed that the Fisher information provides an estimate of the covariance of estimator , i.e.
which could be used to generate a Gaussian estimate of the posterior (see also Tarantola, 2005 for the same result). This work uses the latter expression (Eq. 11) throughout as its familiar form allows for easier interpretation, but inversions using Eq. 10 would yield identical results.
There are alternative compression schemes that could be used. One possibility is to rely entirely on ML models to learn to preserve as much information as possible about the parameters of interest given a set of observations . One approach, for example, aims to maximise the mutual information between model parameters and a learnt set of summary statistics using deep learning (Charnock et al., 2018; Jeffrey et al., 2020). We leave exploring such alternatives to future work.
@@ 3.6 Assessing Posterior Quality
Unfortunately, computing empirical coverage can become very computationally expensive in high-dimensions as many methods rely on explicit density modelling to compute the credible region around the ground truth for quantile estimation (see e.g. Hermans et al., 2021)). Lemos et al. (2023) provided an alternative approach, named Tests of Accuracy with Random Points (TARP), to estimate the coverage much more efficiently. This approach computes the fraction of samples that occur between a random reference point and the true model parameters, allowing one to estimate the observed distribution of credibility levels over a number of inversions and check if they match the expected coverage. A visualisation of this procedure, and how it detects incorrect posteriors, is given in Figs. SSupporting Information–SSupporting Information. This study will utilise TARP to check whether Gaussian-likelihood inversions and SBI produce well-calibrated posterior distributions.
@@ 4 Application to synthetic and real waveform source inversions
In the remainder of this paper we will apply SBI to synthetic and real data earthquake point source inversions for: (i) six independent moment tensor components, fixing the source location; and (ii) space-time location (centroid latitude, longitude, depth and origin time) and the seismic moment tensor (i.e., making a total of 10 source parameters). We will compare SBI with classical Bayesian inversion methods and will carry out synthetic inversions considering both Gaussian and real data noise. Inversion results will be examined in the moment tensor component basis, as well as in a physically interpretable reparametrisation using strike, dip, rake, the non-double-couple component , and moment magnitude . Finally, the SBI approach will be applied to two moderate magnitude earthquakes in the Azores and Madeira regions combining seismic waveforms recorded by stations in the region’s ocean islands and by the recent large-scale UPFLOW ocean bottom seismology experiment in the region (Tsekhmistrenko et al., 2024).
@@ 4.1 Earthquakes Studied and Data Used
The UPFLOW passive seismology experiment deployed 50 and recovered 49 OBSs in a area in the Azores-Madeira-Canary islands region with station spacing km between June 2021 and August 2022. The OBS array is shown in Fig. 4. Most of the instruments had three component broadband seismic sensors and hydrophones, and recorded seismic data with quality as high or even higher than in previous OBS experiments (Tsekhmistrenko et al., 2024). Yet, as expected for OBS data, high noise levels are observed, with noise observations for most stations typically being close to the new high noise model (Tsekhmistrenko et al., 2024). Such high and non-Gaussian noise levels make UPFLOW’s data particularly suitable for testing our SBI approach.
@@ 4.2 Forward Modelling
The time-varying component of the moment tensor describes how the moment couples vary over the duration of the earthquake rupture. This can often play an important role in modelling realistic earthquakes, particularly for large magnitude earthquakes where ruptures last long periods of time. In this study, which focuses on moderate magnitude regional earthquakes, it is assumed that the slip duration is small relative to the shortest period being simulated, i.e. meaning that the slip duration can be neglected. Mathematically it can be treated as a Dirac delta in Eq. 14, i.e. ( is now constant in time), such that the displacement field can be expressed simply as:
@@
Note also that for a fixed source location , the displacement field at each receiver is expressly linear with respect to the moment tensor components. This property is utilised to explore a simplified toy problem in the Section 5.1, since it allows comparison of SBI with analytic results associated with linear Gaussian inference problems.
This study utilises two crustal 1D earth models: one computed beneath station ROSA in the Azores islands (Ferreira et al., 2020), and a Madeira model provided by IPMA. Below the crustal model, LITHO1.0 (Pasyanos et al., 2014) is used until its maximum depth, after which the Preliminary reference Earth model (PREM) 1D (Dziewonski & Anderson, 1981) is used. Green’s functions were computed using AxiSEM-1D (Nissen-Meyer et al., 2016) down to a dominant mesh period of 10 s. Once the databases were computed and processed, Instaseis was used to generate synthetic vertical and horizontal component seismograms at each station. -minute windows starting from one minute before the centroid time are used for each station. Event and noise observations have the instrument response removed, before being processed in exactly the same way as the synthetics. All real and synthetic waveforms are filtered using a 4-corner band-pass filter between and Hz (T s). These relatively long period data are used to mitigate against modelling errors, both due to the simplistic 1D Earth models being used, as well as unmodelled ocean-layer ringing effects observed in the OBS data.
@@ 4.5 Adapting Score Compression to Non-Linear Domains
For such realistic problems that are not expressly linear, selection of the expansion point is key for successful application of optimal score compression. Since the compression is only locally optimal, it stands to reason that accurate inversions require that be located in or near the bulk of the posterior mass. However, this requirement poses a problem in specifying the prior; an “uninformative” prior should cover the entire parameter space, including regions where locally computed breaks down.
@@
This work proceeds by first computing the maximum likelihood solution using a weighted iterative least squares using the approximate Gaussian likelihood. Once the solution for has converged, the relevant quantities and Fisher information can be computed. Finally, the Gaussian posterior covariance estimate , which gives an estimate of the per-parameter standard deviation , can be used to construct a prior over model parameters . In order to avoid an overly informative prior that may bias the final result, instead of using the Gaussian posterior directly, a uniform hypercube centered on is used:

### v2

@@
Earthquake source observations – Waveform inversion – Bayesian inference – Inverse theory – Machine learningThis paper presents a novel framework for full-waveform seismic source inversion using simulation-based inference (SBI). Traditional probabilistic approaches often rely on simplifying assumptions about data errors, which we show can lead to inaccurate uncertainty quantification. SBI addresses this limitation by learning an empirical probabilistic relationship between the parameters and data, without making assumptions about the data errors. This is achieved through the use of specialised machine learning models, known as neural density estimators, which can then be integrated into the Bayesian inference framework. We apply the SBI framework to point-source moment tensor inversions as well as joint moment tensor and time-location inversions. We construct a range of synthetic examples to explore the quality of the SBI solutions, as well as to compare the SBI results with standard Gaussian likelihood-based Bayesian inversions. We then demonstrate that under real seismic noise, common Gaussian likelihood assumptions for treating full-waveform data yield overconfident posterior distributions that underestimate the moment tensor component uncertainties by up to a factor of 3. We contrast this with SBI, which produces well-calibrated posteriors that generally agree with the true seismic source parameters, and offers an order-of-magnitude reduction in the number of simulations required to perform inference compared to standard Monte Carlo sampling techniques. Finally, we apply our methodology to a pair of moderate magnitude earthquakes in the North Atlantic. We utilise seismic waveforms recorded by the recent UPFLOW ocean bottom seismometer array as well as by regional land stations in the Azores, comparing full moment tensor and source-time location posteriors between SBI and a Gaussian likelihood approach. We find that our adaptation of SBI can be directly applied to real earthquake sources to efficiently produce high quality posterior distributions that significantly improve upon Gaussian likelihood approaches.
@@ 1 Introduction
However, there are still some key remaining challenges in seismic source inversions. For example, the study of earthquakes in mid-ocean settings is typically limited by large distances to seismic stations, which provide useful but restricted information about the source process (López-Comino et al., 2015). Local/regional data from marine seismology experiments can help address this issue, but such expensive experiments are usually limited to short time periods and the data collected are often noisy (e.g., Stähler et al., 2016; Tsekhmistrenko et al., 2024). Crucially, robust uncertainty quantification is still lacking in many earthquake source studies, which limits meaningful interpretations (Wéber, 2006). Over the past twenty years, the Bayesian treatment of seismic source inversion has received increasing interest (Lomax et al., 2000; Tarantola, 2005; Vackář et al., 2017; Vasyura-Bathke et al., 2020). Seismic source inversion can be complicated by inter-parameter trade-offs (meaning different parameter combinations fit the data equally well) and strong non-uniqueness (Kawakatsu, 1996; Dufumier & Rivera, 1997; Ford et al., 2010), and previous work has demonstrated that classical statistical analysis of standard errors can significantly underestimate uncertainty in source parameters (Duputel et al., 2012; Valentine & Trampert, 2012). A number of studies have demonstrated that thorough Bayesian treatment of seismic events uncovers uncertainties and parameter trade-offs, often revealing greater complexity than simple maximum likelihood solutions (Duputel et al., 2012; Pugh et al., 2016; Hu et al., 2023). Simutė et al. (2023), for example, demonstrated how inferred focal mechanisms and their distributions are sensitive to the seismic frequencies considered in the inversion, and that a Bayesian analysis of seismic source inversion can produce slightly multimodal posteriors. In addition, recent work has demonstrated that full-waveform probabilistic inversions can be highly sensitive to the noise model parametrisation (Duputel et al., 2012; Mustać & Tkalčić, 2016; Mustać et al., 2018).
@@
To the best of our knowledge this study is the first application of SBI in seismology. We therefore begin this paper by presenting the motivation and prior work in SBI, introducing modern ML techniques for probability distribution modelling, and presenting a data compression technique to treat the full-waveform source inversion problem with SBI. We then present our methodology and apply it to a range of synthetic waveform inversions. Subsequently we use an ocean bottom seismology dataset recently collected in the Atlantic region (the UPFLOW dataset; Ferreira, 2024; Tsekhmistrenko et al., 2024) to study two oceanic earthquakes. This allows us to probe the advantages and pitfalls of using SBI in earthquake source inversions using noisy seismic data.
@@ 3.3 Neural density estimators (NDEs)
The task of inference then relies on finding and training a model capable of modelling complex probability distributions. Much of the success of SBI hinges on the advancements made in the flexibility and efficiency of NDEs over the last decade. Neural density estimation is a conceptually simple task at its core; given a dataset of observations of where runs over all examples in the dataset, a NDE must be trained to estimate the density across . Conditional probabilities can be similarly treated, with a dataset of allowing the empirical modelling of e.g. density . In practice, the challenge in performing this modelling is to design a flexible density estimator with a tractable (ideally easy to compute) density function.
@@
where is the Jacobian of T, i.e., the matrix of partial derivatives with respect to components of . One type of normalising flow are autoregressive models, where is decomposed into its elements . One can then proceed by learning to model a succession of conditional distributions, thereby constructing the full probability distribution by applying the chain rule:
@@
These flows also have the useful property that they can be composed together, in effect learning to parametrise successive transformations, allowing for significant improvements in the flexibility of the model. MAFs, which we utilise in this work, are constructed by stacking together several MADE networks, indexed by , and varying the autoregression order from Eq. 5 between each MADE. In this work, MAFs are trained to model the posterior , so each transformation parametrised by a MADE is also conditioned on the observation . This architecture is depicted in Fig. 2.
Once the network architecture has been selected, the NDE can be trained to model the observed density . The mismatch between the distributions can be quantified using the Kullback–Leibler (KL) divergence, leading to the following objective function:
where the first term can be ignored as it is a constant that does not depend on the network parameters (see e.g., Papamakarios et al., 2021). For a dataset , posterior estimation can therefore be achieved through the simple objective , i.e. the network should aim to maximise the probability density it outputs over the observed data. This straightforward formulation can be combined with standard stochastic gradient descent techniques common to most ML applications to train the NDE.
@@ 3.5 Optimal Score Compression
where the Fisher information matrix is defined as for log-likelihood , and . Here, represents an expectation over the model parameters. Eq. 8 states that the variance of the statistic over the model parameters can be bounded from below (where equality holds for an “optimal” statistic). Stated informally, to maximally constrain model parameters when inferring them from an observation we should select a compression scheme such that the variance over is minimised for a given .
@@
where is the synthetic data at the expansion point , and is the gradient of this data vector with respect to the model parameters. Eq. 11 thus defines the compression operator for the optimal score compression scheme.
@@
where the Fisher information for the Gaussian likelihood in Eq. 9. This form of the summary statistic corresponds to a local maximum likelihood estimate (MLE) of the model parameters, , given the observation ; equivalently, it is an estimate of the model parameters after a single step of least-squares from expansion point . This work uses the latter expression (Eq. 12) throughout as its familiar form allows for easier interpretation, but inversions using Eq. 11 would yield identical results.
There are alternative compression schemes that could be used. One possibility is to rely on ML models to learn to preserve as much information as possible about the parameters of interest given a set of observations . This is an active area of research with a wide range of strategies for training deep neural networks to compress high dimensional observations (Fluri et al., 2021; Lu et al., 2022; Sharma et al., 2024), see e.g. Lanzieri et al. (2024) for a review. One approach, for example, aims to maximise the mutual information between model parameters and a learnt set of summary statistics using deep learning (Charnock et al., 2018; Jeffrey et al., 2020; Prelogović & Mesinger, 2024). We leave exploring such alternatives for seismic source inversion to future work.
@@ 3.6 Assessing Posterior Quality
Unfortunately, computing empirical coverage can become very computationally expensive in high-dimensions as many methods rely on explicit density modelling to compute the credible region around the ground truth for quantile estimation (see e.g. Hermans et al., 2021)). Lemos et al. (2023) provided an alternative approach, named Tests of Accuracy with Random Points (TARP), to estimate the coverage much more efficiently. This approach computes the fraction of samples that occur between a random reference point and the true model parameters, allowing one to estimate the observed distribution of credibility levels over a number of inversions and check if they match the expected coverage. A visualisation of this procedure, and how it detects incorrect posteriors, is given in Figs SSupporting Information–SSupporting Information. This study will utilise TARP to check whether Gaussian-likelihood inversions and SBI produce well-calibrated posterior distributions.
@@ 4 Application to synthetic and real waveform source inversions
In the remainder of this paper we will apply SBI to synthetic and real data earthquake point source inversions for: (i) six independent moment tensor components, fixing the source location; and (ii) space-time location (centroid latitude, longitude, depth and origin time) and the seismic moment tensor (i.e., making a total of 10 source parameters). We will compare SBI with classical Bayesian inversion methods and will carry out synthetic inversions considering both Gaussian and real data noise. Inversion results will be examined in the moment tensor component basis, as well as in a physically interpretable reparametrisation using strike, dip, rake, the non-double-couple component (following the convention from Giardini, 1984), and moment magnitude . Finally, the SBI approach will be applied to two moderate magnitude earthquakes in the Azores and Madeira regions combining seismic waveforms recorded by stations in the region’s ocean islands and by the recent large-scale UPFLOW ocean bottom seismology experiment in the region (Tsekhmistrenko et al., 2024).
@@ 4.1 Earthquakes Studied and Data Used
The UPFLOW passive seismology experiment deployed 50 and recovered 49 OBSs in a area in the Azores-Madeira-Canary islands region with station spacing km between June 2021 and August 2022. The OBS array is shown in Fig. 4. Most of the instruments had three component broadband seismic sensors and hydrophones, and recorded seismic data with quality as high or even higher than in previous OBS experiments (Tsekhmistrenko et al., 2024). Yet, as expected for OBS data, high noise levels are observed, with noise observations for most stations typically being close to the new high noise model (Tsekhmistrenko et al., 2024). Such high and non-Gaussian noise levels make UPFLOW’s data particularly suitable for testing our SBI approach. All UPFLOW data have clock drift estimated and corrected for (Tsekhmistrenko et al., 2024; Cabieces et al., 2024). While future work will use data corrected for tilt and compliance, these corrections should not affect this study’s results because the data’s dominant period is s. This is at the limit for which tilt and compliance corrections become substantial for the UPFLOW dataset (Tsekhmistrenko et al., 2024).
@@ 4.2 Forward Modelling
The time-varying component of the moment tensor describes how the moment couples vary over the duration of the earthquake rupture. This can often play an important role in modelling realistic earthquakes, particularly for large magnitude earthquakes where ruptures last long periods of time. In this study, which focuses on moderate magnitude regional earthquakes, it is assumed that the slip duration is small relative to the shortest period being simulated, i.e. meaning that the slip duration can be neglected. Mathematically it can be treated as a Dirac delta in Eq. 14, i.e. , where is now constant in time, such that the displacement field can be expressed simply as:
@@
Note also that for a fixed source location , the displacement field at each receiver is expressly linear with respect to the moment tensor components. This property is utilised to explore a simplified toy problem in the Section 5.1, since it allows comparison of SBI with analytic results associated with linear Gaussian inference problems (Tarantola, 2005; Alsing & Wandelt, 2018).
This study utilises two crustal 1D earth models: one computed beneath station ROSA in the Azores islands (Ferreira et al., 2020), and a Madeira model provided by IPMA. Below the crustal model, LITHO1.0 (Pasyanos et al., 2014) is used until its maximum depth, after which the Preliminary reference Earth model (PREM) 1D (Dziewonski & Anderson, 1981) is used. Green’s functions were computed using AxiSEM-1D (Nissen-Meyer et al., 2016) down to a dominant mesh period of 10 s. Once the databases were computed and processed, Instaseis was used to generate synthetic vertical and horizontal component seismograms at each station. -minute windows starting from one minute before the centroid time are used for each station. Event and noise observations have the instrument response removed, before being processed in exactly the same way as the synthetics. All real and synthetic waveforms are filtered using a 4-corner band-pass filter between and Hz (T s).
These relatively long period data are used to mitigate against modelling errors, both due to the simplistic 1D Earth models being used, as well as unmodelled ocean-layer ringing effects observed in the OBS data. In our setting, particularly when treating waveforms with high signal-to-noise ratios, we expect large modelling errors that dominate contributions to the uncertainty and bias the inference results. This is relevant for the moderate magnitude events studied in this manuscript, and we present a discussion of this issue in Section 6.1.2.
@@ 4.5 Adapting Score Compression to Non-Linear Domains
For such realistic problems that are not expressly linear, selection of the expansion point is key for successful application of optimal score compression. Since the compression is only locally optimal, it stands to reason that accurate inversions require that be located in or near the bulk of the posterior mass. However, this requirement poses a problem in specifying the prior; an “uninformative” prior should cover the entire parameter space, including regions where locally computed gradients break down.
@@
This work proceeds by first computing the maximum likelihood solution using a weighted iterative least squares using the approximate Gaussian likelihood. Once the solution for has converged, the relevant quantities and Fisher information can be computed. Finally, the Gaussian posterior covariance estimate (see e.g., Tarantola, 2005), which gives an estimate of the per-parameter standard deviation , can be used to construct a prior over model parameters . In order to avoid an overly informative prior that may bias the final result, instead of using the Gaussian posterior directly, a uniform hypercube centered on is used: